OpenStudy (firejay5):

Solve each system of equations by using substitution: 13. 2j - 3k = 3 j + k = 14 14. 2r + s = 11 6r - 2s = -2

OpenStudy (anonymous):

First find at least one constant variable to substitute, for problem 13, I would leave either k or j by its self in j+k=14, so than you would have J=14-k, since you have J, substitute it with the top equation, 2(14-k)-3k=3. I presume you can go from here.

OpenStudy (anonymous):

Do you see how the steps progresses>?

OpenStudy (firejay5):

yes

OpenStudy (anonymous):

Awesome, YOU GOT THIS MAN.

OpenStudy (firejay5):

IKR!!! :D

OpenStudy (firejay5):

@rendition91 Do number 14

OpenStudy (anonymous):

How about I point you in the right direction? What do you think you would do first.

OpenStudy (anonymous):

First you need to leave one variable like r or s by it's self in order to substitute. From those two equation in number 14, Which one seems simpler to define one variable.

OpenStudy (firejay5):

@rendition91 How would you solve the second equation on #14???

OpenStudy (anonymous):

Same way you would on the first, Okay Here is my only help that I am going to give you ha sorry I just want you to understand the concept so you can progress on your own. Just like the first, single out one of the equation to a single variable to use as substitution. It looks ideal to single out one of the variables in 2r+s=11, substract that 2r and you would have s=11-2r, Now you have a single variable to substitute with the second equation. Implement that s with 6r-2s=-2, Therefore you would have 6r-2(11-2r)=-2, Now it is your turn to complete it.